OpenStudy (anonymous):

Simplify (x + 3)(x2 + 4x – 2)

OpenStudy (anonymous):

What do you mean simplify?

OpenStudy (anonymous):

I'm not quite sure, I'm terrible at Algebra. I believe it uses the FOIL method, but I'm not 100% on that..

OpenStudy (anonymous):

Yeah that makes more sense to foil them. So do you want I foil them for you?

OpenStudy (anonymous):

If you can please, I wasn't in school whene we learned the FOIL method.

OpenStudy (anonymous):

= x^3+4x^2-2x+3x^2+12x-6 =x^3+7x^2+10x-6

jimthompson5910 (jim_thompson5910):

Hint (x + 3)(x^2 + 4x – 2) is the same as x(x^2 + 4x – 2) + 3(x^2 + 4x – 2)